OpenStudy (anonymous):

Does the equation x^2-9 represent the equation (x-3)^2

OpenStudy (anonymous):

\[x^2-9=(x-3)^2\] \[x^2-9=x^2-6x+9\] no

OpenStudy (anonymous):

wow I was way off, thanks @girlscum

OpenStudy (anonymous):

no problem!!